OpenStudy (timaashorty):

If you're happy and you know it... AYEE you viewed this. Might as well give me a Medal! (; Yes or yes ? you have no other choice

OpenStudy (timaashorty):

Answer choices:

OpenStudy (shamil98):

universe?

OpenStudy (anonymous):

b

OpenStudy (timaashorty):

Lmaoo *Inverse

OpenStudy (shamil98):

oh that makes more sense

OpenStudy (timaashorty):

Positive batman ? ;P

OpenStudy (anonymous):

b for batman trust me

OpenStudy (shamil98):

y=7x^2-3 solve for x. y + 3 = 7x^2 y + 3 / 7 = x^2 √(y+3)/7) = x yes hes right

OpenStudy (anonymous):

i looked at the function and did the algebra in yocto seconds in my head so yeah im positive!

OpenStudy (anonymous):

yocto seconds being = 10^-35 seconds fyi

OpenStudy (timaashorty):

Sorry Sham I had to give him that medal lol thanks both of you(:

OpenStudy (ranga):

fyi: yocto second is 10^-24 second.

OpenStudy (anonymous):

google faster

OpenStudy (anonymous):

err slower

OpenStudy (anonymous):

and you're still wrong!!!!!